# This Version: March 9, 2024

### v1

@@ 4 Main Results
The state, action, and parameter spaces are lattices, and for any primitive $p\in P,$ $u(s,x,p)$ is supermodular in $(s,x)$ and increasing in $s.$
Our first assumption is rather standard and emphasises a lattice structure on the state, action, and parameter spaces. As evinced by the examples in this paper and in general, for many economic applications, the state, action, and parameter spaces are a subset of the real line $\mathbb{R},$ for which the lattice assumption is trivially satisfied in the natural order. Further, we assume that the payoff function is supermodular in state and action and is increasing in the state variable. A payoff function $u:\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}$ is supermodular if, for all $(s_{1},x_{1}),(s_{2},x_{2})\in\mathbb{S}\times\mathbb{X},u((s_{1},x_{1})\vee(s_{2},x_{2}))+u((s_{1},x_{1})\wedge(s_{2},x_{2}))\geq u(s_{1},x_{1})+u(s_{2},x_{2}).$ This implies that states and actions exhibit complementarity in the sense that marginal contribution to the payoff of increasing the action increases with a higher state. The supermodularity in states and actions and monotononicity (in state) of the payoff function alongwith our next assumption guarantees that optimal policy correspondence is increasing in the state variable, for any given model distribution $\mu,$ and primitive $p\in P.$
@@
Notice that this assumption is exclusively on the set of models. It requires the model transition functions to be stochastically increasing and stochastically supermodular in states and actions. This implies that for every model $Q_{\theta},$ a higher current state and action increases the probability of observing a higher state in the next period and that an incremental amount of action increases this probability, the lower the state is in the current period.373737 For example, consider the following AR(1) process, $s_{t+1}=\theta s_{t}+\epsilon_{t+1},$ where $\epsilon_{t+1}$ is distributed normally with mean 0 and variance $\sigma^{2}.$ Then for every $\theta,$ $Q_{\theta}$ is stochastically increasing. See Examples 2 and 3 for illustrations on stochastic supermodularity. Assumptions 1 and 2 are consistent with the structure one requires for a correctly specified environment and the reader is referred to Theorem 3.9.2 in Topkis (1998) for a similar set of assumptions. For a given model distribution $\mu,$ and primitive $p,$ the above two assumptions guarantee an increasing optimal policy correspondence $G(s,\mu,p).$
@@
For any given $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}),$ a SMDP $(Q,Q_{\Theta})$ is point-identified if $\theta,\theta^{\prime}\in\Theta(m;Q)$ implies that $\theta=\theta^{{}^{\prime}}.$ It is point-identified if it is point-identified for all $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}).$ [^38] The reader is referred here to Lewbel (2019) for an exposition on identification problem in economics.
The identification assumption is satisfied for many applications in the existing misspecification literature and also plays an important role in the convergence of the steady state behavior of an agent in a SMDP to the Berk-Nash equilibrium; see examples in Esponda and Pouzo (2016), EP, and, Esponda, Pouzo, and Yamamoto (2021). For our comparative statics purposes, identification ensures that we always get a unique selection $\theta_{Q}(m)$ from the best-fit set $\Theta(m;Q).$ A function $\theta_{Q}(m),$ maps a distribution $m$ into a real number in $\Theta$ is said to be increasing in the usual stochastic order $\succsim_{\mathrm{st}}$, if $\theta_{Q}(m_{2})\geq\theta_{Q}(m_{1})$ whenever $m_{2}\succsim_{\text{st }}m_{1}.$ For our last assumption, we require the weighted Kullback-Liebler divergence to exhibit the single-crossing property in the parameter $\theta$ and distribution $m$.
@@
$K_{Q}(\theta;m)$ satisfies the single crossing property in $(\theta;m),$ that is, for $\theta_{1}\leq\theta_{2},$ and for $m_{2}\succsim_{st}m_{1},$ $\delta(m_{1})=K_{Q}(\theta_{2},m_{1})-K_{Q}(\theta_{1},m_{1})\geq(<)0$ implies $\delta(m_{2})=K_{Q}(\theta_{2},m_{2})-K_{Q}(\theta_{1},m_{2})\geq(<)0.$
@@
Sufficient Condition 1: For any two models $\theta_{1},\theta_{2}\in\Theta,$ such that $\theta_{1}<\theta_{2},$ define the log-likelihood ratio, $L(s^{\prime}|s,x)=\ln(D_{\theta_{2}}(s^{\prime}|s,x))-\ln(D_{\theta_{1}}(s^{\prime}|s,x)),$ where $D_{\theta}(s^{\prime}|s,x)$ are the Radon-Nikodym derivatives, introduced for measuring KL divergences in (5). Then the models are said to follow the generalized likelihood ratio property if the expectation of $L$ with respect to the true distribution $Q(\cdot|s,x)$ is increasing in state and actions, $(s,x).$ [^40] Ross (1987) contains an useful discussion regarding the existence of the expectation of the log-likelihood ratio statistic. For instance, if the true, and the model distributions belong to the Gaussian family, this statistic would always exist.
@@
EP and ADGK provide proofs for the existence of a Berk-Nash equilibrium in an SMDP with finite and infinite environments, respectively. While the former proves the theorem for finite states and actions, ADGK takes the finite result of EP as given and extends it to more naturally appealing instances of infinite environments, using novel tools in non-standard analysis. In contrast, our existence result relies on the assumed monotonicity and identification properties of the structure of our problem. It is important to mention here that we do rely on the regular SMDP definition in ADGK, most importantly on part (iv) of Definition 3, because it allows us to consider distributions that have unbounded Radon-Nikodym derivatives. This is particularly relevant when dealing with distributions over infinite spaces, such as the AR(1) process. [^41] Although our examples are based on unbounded state spaces, our results in this paper are established for compact state and action spaces. For unbounded state spaces, one continues to rely on ADGK’s structure and existence results. Further, the set of Berk-Nash equilibria will be compact and in particular, there will always exist the least and the greatest Berk-Nash equilibrium, $m^{*}$ and corresponding beliefs $\mu^{*}$ over the best-fit inferred models. [^42] The argument for the least and greatest follows from Theorem 4 in Acemoglu and Jensen (2015) for the non-lattice case. Also, see Footnote 9 (pp. 1389) in Hopenhayn and Prescott (1992) for several instances of compact subsets of measures in economic problems.
@@
For any given belief $\mu,$ a change in a primitive of the SMDP from $p_{1}$ to $p_{2}$ is a positive (negative) shock if $G(s,\mu,p_{2})$ is greater than $G(s,\mu,p_{1})$ in the strong set order. That is, for all $y_{1}\in G\left(s,\mu,p_{1}\right)$ and $y_{2}\in G\left(s,\mu,p_{2}\right),$ the join $y_{1}\vee y_{2}\in G\left(s,\mu,p_{2}\right)$ and the meet $y_{1}\wedge y_{2}\in G\left(s,\mu,p_{1}\right).$ Further, for a given primitive $p\in P$, a change in beliefs over models from $\mu_{1}$ to $\mu_{2}$ is a positive (negative) shock if $G(s,\mu,p)$ is ascending in $\mu$ from $\mu_{1}$ to $\mu_{2},$ that is, if $y_{1}\vee y_{2}\in G(s,\mu_{2},p)$ and $y_{1}\wedge y_{2}\in G(s,\mu_{1},p)$ for all $y_{1}\in G(s,\mu_{1},p)$ and $y_{2}\in G(s,\mu_{2},p).$
@@
Our framework is more general to that of Heidhues, Kőszegi, and Strack (2018), and therefore, the sharp predictions that they have regarding “perverse” comparative statics for a stylized production problem with a loss function is something that we are not able to explain here.
@@
Short proof sketch: The proof of Theorem 2 follows a three-step proof structure as in Acemoglu and Jensen (2015). The first step involves showing that under Assumptions 1 and 2, and for any fixed model distribution $\mu$, the set of Berk-Nash equilibrium $m$ obtained via the fixed points of the equilibrium mapping $T$ will be Type I and Type II increasing in the primitives $p.$ Using the set of fixed points in the first step, the second step involves constructing a mapping $\hat{\theta}$ that for any model distribution $\mu$ and primitive $p,$ gives a set of model distributions. This mapping is constructed from part (b) of Definition 5 wherein for each of the Berk-Nash equilibrium $m$ obtained in step 1, the construction gives a set of $\mu^{\prime}s.$ It is the fixed points of this map that are our equilibrium distribution given $p$. The third and final step shows that if the least and the greatest selections of the mapping $\hat{\theta}$ are increasing in primitives, then the associated fixed points are increasing in $p$.
@@
Acemoglu and Jensen (2015) gives several sufficient conditions to identify positive shocks for their environment. All of their results (Lemmas 1-3) translate to our case, [^44] The reader is referred to pp. 604-605 in Acemoglu and Jensen (2015). albeit with some moderation for the endogenous dynamic programs. For example, a change in the primitive $p$ that influences the decision problem through the utility function, such as the risk-aversion parameter, is a positive shock if the utility function $u(s,x,p)$ has increasing differences in $x$ and $p.$ Similarly, under Assumptions 1 and 2, a change in discount factor $\beta$ and model distribution $\mu$ is a positive shock.
@@
Suppose the hypothesis in Theorem 1 continue to hold. If a change in beliefs over models is a positive shock, then an increase in the parameter set under the strong set order leads to an increase in the least and the greatest equilibrium best-fit models.
Our last assumption modifies the single crossing property for instances when distributions over states and actions are ordered in the increasing and convex stochastic order.
@@
Sufficient Condition 2: For any two models $\theta_{1},\theta_{2}\in\Theta,$ such that $\theta_{1}<\theta_{2},$ define the log-likelihood ratio, $L(s^{\prime}|s,x)=\ln(D_{\theta_{2}}(s^{\prime}|s,x))-\ln(D_{\theta_{1}}(s^{\prime}|s,x)).$ Then the models are said to follow the expected likelihood ratio property in increasing and convex order if the expectation of $L$ with respect to the true distribution $Q$ is increasing and convex in state and actions, $(s,x).$ Our final theorem addresses the monotonicity of Berk-Nash equilibrium for the increasing and convex order.
@@
Suppose assumptions 1-3 and assumption 5 holds. Then a positive (negative) shock to the primitives will lead to an increase in the least and greatest Berk-Nash equilibrium in the increasing convex order if increases in beliefs over models are positive (negative) shocks.

### v2

@@ 4 Main Results
The state and action spaces are lattices, and for any primitive $p\in P,$ $u(s,s^{\prime},x,p)$ is supermodular in $(s,x)$ and increasing in current state $s$ and future state $s^{\prime}.$
Our first assumption is rather standard and emphasises a lattice structure on the state, action, and parameter spaces. As evinced by the examples in this paper and in general, for many economic applications, the state, action, and parameter spaces are a subset of the real line $\mathbb{R},$ for which the lattice assumption is trivially satisfied in the natural order. Further, we assume that the payoff function is supermodular in current state and action and is increasing in the future state variable. A payoff function $u:\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}$ is supermodular if, for all $(s_{1},x_{1}),(s_{2},x_{2})\in\mathbb{S}\times\mathbb{X},u((s_{1},x_{1})\vee(s_{2},x_{2}))+u((s_{1},x_{1})\wedge(s_{2},x_{2}))\geq u(s_{1},x_{1})+u(s_{2},x_{2}).$ This implies that states and actions exhibit complementarity in the sense that marginal contribution to the payoff of increasing the action increases with a higher state. The supermodularity in states and actions and monotononicity (in state) of the payoff function alongwith our next assumption guarantees that optimal policy correspondence is increasing in the state variable, for any given model distribution $\mu,$ and primitive $p\in P.$
@@
Notice that this assumption is exclusively on the set of models. It requires the model transition functions to be stochastically increasing and stochastically supermodular in states and actions. This implies that for every model $Q_{\theta},$ a higher current state and action increases the probability of observing a higher state in the next period and that an incremental amount of action increases this probability, the higher the state is in the current period.393939 For example, consider the following AR(1) process, $s_{t+1}=\theta s_{t}+\epsilon_{t+1},$ where $\epsilon_{t+1}$ is distributed normally with mean 0 and variance $\sigma^{2}.$ Then for every $\theta,$ $Q_{\theta}$ is stochastically increasing. See Examples 2 and 3 for illustrations on stochastic supermodularity. Assumptions 1 and 2 are consistent with the structure one requires for a correctly specified environment and the reader is referred to Theorem 3.9.2 in Topkis (1998) for a similar set of assumptions. For a given model distribution $\mu,$ and primitive $p,$ the above two assumptions guarantee an increasing optimal policy correspondence $G(s,\mu,p).$
@@
For any given $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})$, we assume that the SMDP $(Q,Q_{\Theta})$ is point-identified, meaning that $\theta,\theta^{\prime}\in\Theta(m;Q)$ implies $\theta=\theta^{\prime}$. Moreover, we assume this holds for all $m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})$. [^40] The reader is referred here to Lewbel (2019) for an exposition on identification problem in economics.
The identification assumption is satisfied for many applications in the existing misspecification literature and also plays an important role in the convergence of the steady state behavior of an agent in a SMDP to the Berk-Nash equilibrium; see examples in Esponda and Pouzo (2016), EP, and, Esponda, Pouzo, and Yamamoto (2021). For our comparative statics purposes, identification ensures that we always get a unique selection $\theta_{Q}(m)$ from the best-fit set $\Theta(m;Q).$ A function $\theta_{Q}(m),$ maps a distribution $m$ into a real number in $\Theta$ is said to be increasing in the usual stochastic order $\succsim_{\mathrm{st}}$, if $\theta_{Q}(m_{2})\geq\theta_{Q}(m_{1})$ whenever $m_{2}\succsim_{\text{st }}m_{1}.$ For our last assumption, we require the weighted Kullback-Leibler divergence to exhibit the single-crossing property in the parameter $\theta$ and distribution $m$.
@@
$K_{Q}(\theta;m)$ satisfies the single crossing property in $(\theta;m),$ that is, for $\theta_{1}\leq\theta_{2},$ and for $m_{2}\succsim_{st}m_{1},$ $\delta(m_{1})=K_{Q}(\theta_{2},m_{1})-K_{Q}(\theta_{1},m_{1})\geq(<)0$ implies $\delta(m_{2})=K_{Q}(\theta_{2},m_{2})-K_{Q}(\theta_{1},m_{2})\geq(<)0.$ [^41] It is important to note that as long as the divergence “measure” satisfies a form of single crossing, as outlined in Assumptions 4 and 5, the comparative statics results remain valid. In this regard, the KL divergence serves more as an example than a critical feature for monotonicity. However, our emphasis on the KL divergence stems from the fact that the Berk-Nash solution concept identifies it as a potential limit of the misspecified learning process. Alternatively, if another divergence measure were to characterize the limit, it would need to satisfy the single crossing property for the comparative statics results to hold. I thank a referee for this observation.
@@
Sufficient Condition 1: For any two models $\theta_{1},\theta_{2}\in\Theta,$ such that $\theta_{1}<\theta_{2},$ define the log-likelihood ratio, $L(s^{\prime}|s,x)=\ln(D_{\theta_{2}}(s^{\prime}|s,x))-\ln(D_{\theta_{1}}(s^{\prime}|s,x)),$ where $D_{\theta}(s^{\prime}|s,x)$ are the Radon-Nikodym derivatives, introduced for measuring KL divergences in (5). Then the models are said to follow the expected likelihood ratio property if the expectation of $L$ with respect to the true distribution $Q(\cdot|s,x)$ is increasing in both state and action, $(s,x).$ [^43] Ross (1987) contains a useful discussion regarding the existence of the expectation of the log-likelihood ratio statistic. For instance, if the true, and the model distributions belong to the Gaussian family, this statistic would always exist.
@@
EP and ADGK provide proofs for the existence of a Berk-Nash equilibrium in an SMDP with finite and infinite environments, respectively. While the former proves the theorem for finite states and actions, ADGK takes the finite result of EP as given and extends it to more naturally appealing instances of infinite environments (compact-metric spaces), using novel tools in non-standard analysis. In contrast, our proof of existence result relies on the assumed monotonicity and identification properties of the structure of our problem, given we are housed in the Euclidean space. It is important to mention here that we do rely on the regular SMDP definition in ADGK, most importantly on part (iv) of Definition 3, because it allows us to consider distributions that have unbounded Radon-Nikodym derivatives. This is particularly relevant when dealing with distributions over infinite spaces, such as the AR(1) process. [^44] While our examples focus on unbounded state spaces, the results in this paper apply to compact state and action spaces, which do not directly impact the comparative statics properties. For unbounded state spaces, one continues to rely on ADGK’s structure and existence results. Further, the set of Berk-Nash equilibria will be compact and in particular, there will always exist the least and the greatest Berk-Nash equilibrium, $m^{*}$ and corresponding beliefs $\mu^{*}$ over the best-fit inferred models. [^45] The argument for the least and greatest follows from Theorem 4 in Acemoglu and Jensen (2015) for the non-lattice case. Also, see Footnote 9 (pp. 1389) in Hopenhayn and Prescott (1992) for several instances of compact subsets of measures in economic problems.
@@
For any given belief $\mu,$ a change in a primitive of the SMDP from $p_{1}$ to $p_{2}$ is a positive (negative) shock if $G(s,\mu,p_{2})$ is greater than $G(s,\mu,p_{1})$ in the strong set order. That is, for all $y_{1}\in G\left(s,\mu,p_{1}\right)$ and $y_{2}\in G\left(s,\mu,p_{2}\right),$ the join $y_{1}\vee y_{2}\in G\left(s,\mu,p_{2}\right)$ and the meet $y_{1}\wedge y_{2}\in G\left(s,\mu,p_{1}\right).$ Further, for a given primitive $p\in P$, a change in beliefs over models in the usual order stochastic dominance sense from $\mu_{1}$ to $\mu_{2}$ is a positive (negative) shock if $G(s,\mu,p)$ is ascending in $\mu$ from $\mu_{1}$ to $\mu_{2},$ that is, if $y_{1}\vee y_{2}\in G(s,\mu_{2},p)$ and $y_{1}\wedge y_{2}\in G(s,\mu_{1},p)$ for all $y_{1}\in G(s,\mu_{1},p)$ and $y_{2}\in G(s,\mu_{2},p).$
@@
Heidhues, Kőszegi, and Strack (2018), henceforth HKS, identify an interesting result of self-defeating learning in misspecified environments. They examine how a myopic agent, with overconfident beliefs—unrealistically high expectations—about their ability, engages in a learning process that leads them away from optimal actions. Technically, the HKS framework is not directly applicable here, as it is not a Markov Decision Process (MDP) per se. However, we adapt its basic structure to formulate it as an MDP, demonstrating how the key assumptions in HKS align with ours, and deriving the comparative statics results accordingly. [^47] I am deeply grateful to an anonymous referee for urging me to make this connection.
In each period $t\in\{1,2,3,\ldots\}$, the agent produces an observable output according to the law of motion:
$q_{t+1}=Q\bigl{(}x_{t},a,\Phi\bigr{)}+\varepsilon_{t+1},$
where $x_{t}\in[\underline{x},\bar{x}]$ is the action chosen by the agent at time $t$, $a\in\mathbb{R}$ represents the agent’s constant ability, and $\Phi\in[\underline{\phi},\bar{\phi}]$ is a fixed but unobservable fundamental parameter. The term $\varepsilon_{t+1}$ represents random noise, which is assumed to be independent and identically distributed (i.i.d.), following the assumptions specified in HKS. The agent is assumed to be overoptimistic about his ability: although his true ability is $A$, he believes his ability is $\tilde{a}$, where $\tilde{a}>A$. We define the degree of overconfidence as $\Delta=|\tilde{a}-A|$. Given this, the agent updates his beliefs about the fundamental parameter $\Phi$ in a Bayesian manner. He discounts future outcomes with a factor $\delta\in(0,1)$, and in each period, chooses a current action $x_{t}$ along with a strategy for future actions, contingent on the history of the game and his learning of the fundamental, in order to maximize his discounted expected output. The agent’s objective is thus to solve the following dynamic optimization problem:
$\max_{x_{t},x_{t+1},\dots}\mathbb{E}_{t}\left[\sum_{t=0}^{\infty}\delta^{t}\,q_{t+1}\right],$
where the expectation is taken with respect to the agent’s beliefs about the future evolution of the state $\Phi$ and the corresponding output $q_{t+1}$, and subject to the law of motion given by the equation above.
The steady state of this dynamic optimization problem is characterized by a Berk-Nash equilibrium, which consists of a steady-state distribution over the state variable and a corresponding action choice $x^{*}$, together with an equilibrium belief $\Phi^{*}$ about the fundamental. In equilibrium, the agent’s Bellman equation satisfies
$V\bigl{(}q_{t},a,\Phi^{*}\bigr{)}=\max_{x^{*}\in[\underline{x},\bar{x}]}\;\mathbb{E}\Bigl{[}\,q_{t+1}+\delta\,V\bigl{(}q_{t+1},a,\Phi^{*}\bigr{)}\Bigr{]},$ (14)
where the expectation is taken over the distribution of $\varepsilon_{t+1}$. Furthermore, the equilibrium belief $\Phi^{*}$ must be consistent with the agent’s subjective model. In particular, $\Phi^{*}$ satisfies the condition
$Q\bigl{(}x^{*},A,\Phi\bigr{)}=Q\bigl{(}x^{*},\tilde{a},\Phi^{*}\bigr{)},$ (15)
which arises as the solution to the problem of minimizing the Kullback-Leibler (KL) divergence between the true and the subjective distributions. [^48] The reader is referred to Lemma 7, page 1190 in HKS.
We prove that in the non-myopic misspecified MDP version of HKS, higher overconfidence results in lower equilibrium effort. This follows from the assumptions in HKS that correspond to our Assumptions 1-4. First, HKS’s Assumption 1, $Q_{ax}\leq 0$, translates in our framework to a change in the primitive (ability) that qualifies as a negative shock due to submodularity. Second, the optimal effort level is always monotonic (increasing) in the change in fundamental, which translates to a change in beliefs about the fundamental in a usual-order stochastic sense, thus a positive shock. Third, for any effort level, there is a unique fundamental consistent with the output produced. This follows from $Q_{\phi}>0$ in HKS, satisfying Assumption 3 in our paper. Finally, since $Q_{x\phi}>0$ in HKS, this implies that optimal action and the fundamental are monotonic in each other. [^49] We assume, without loss of generality, that the noise $\epsilon_{t+1}$ follows a normal distribution and, therefore, satisfies regularity. Readers are referred to page 1163 in HKS for a listing of these assumptions.
@@
Short proof sketch: The proof of Theorem 2 follows a three-step proof structure as in Acemoglu and Jensen (2015). The first step involves showing that under Assumptions 1 and 2, and for any fixed model distribution $\mu$, the set of Berk-Nash equilibriua $m$ obtained via the fixed points of the equilibrium mapping $T$ will be Type I and Type II increasing in the primitives $p.$ Using the set of fixed points in the first step, the second step involves constructing a mapping $\hat{\theta}$ that for any model distribution $\mu$ and primitive $p,$ gives a set of model distributions. This mapping is constructed from part (b) of Definition 5 wherein for each of the Berk-Nash equilibrium $m$ obtained in step 1, the construction gives a set of $\mu^{\prime}s.$ It is the fixed points of this map that are our equilibrium distribution given $p$. The third and final step shows that if the least and the greatest selections of the mapping $\hat{\theta}$ are increasing in primitives, then the associated fixed points are increasing in $p$.
@@
Acemoglu and Jensen (2015) gives several sufficient conditions to identify positive shocks for their environment. All of their results (Lemmas 1-3) translate to our case, [^50] The reader is referred to pp. 604-605 in Acemoglu and Jensen (2015). albeit with some moderation for the endogenous dynamic programs. For example, a change in the primitive $p$ that influences the decision problem through the utility function, such as the risk-aversion parameter, is a positive shock if the utility function $u(s,x,p)$ has increasing differences in $x$ and $p.$ Similarly, under Assumptions 1 and 2, an increase in discount factor $\beta$ is a positive shock.
@@
Suppose the hypothesis in Theorem 1 continue to hold. If a change in beliefs over models in the usual stochastic order is a positive shock, then an increase in the parameter set under the strong set order leads to an increase in the least and the greatest equilibrium best-fit models.
Our last assumption modifies the single crossing property for instances when distributions over states and actions are ordered in the increasing and convex stochastic order. Our final theorem addresses the monotonicity of Berk-Nash equilibrium for the increasing and convex order.
@@
Sufficient Condition 2: For any two models $\theta_{1},\theta_{2}\in\Theta,$ such that $\theta_{1}<\theta_{2},$ define the log-likelihood ratio, $L(s^{\prime}|s,x)=\ln(D_{\theta_{2}}(s^{\prime}|s,x))-\ln(D_{\theta_{1}}(s^{\prime}|s,x)).$ Then the models are said to follow the expected likelihood ratio property in increasing and convex order if the expectation of $L$ with respect to the true distribution $Q$ is increasing and convex in state and actions, $(s,x).$
@@
Suppose assumptions 1-3 and assumption 5 holds. Then a positive (negative) shock to the primitives will lead to an increase in the least and greatest Berk-Nash equilibrium in the increasing convex order if change in beliefs in the usual stochastic order sense over models are positive (negative) shocks.